# End-to-End and Highly-Efficient Differentiable Simulation for Robotics

### v1

@@ Acknowledgments
Appendix C - Collision detection contribution
Given one contact frame $c$ between body $1$ and body $2$, the contact Jacobian is
@@
where $E=\begin{pmatrix}\text{Id}_{3}&0_{3,3}\\ 0_{3,3}&0_{3,3}\end{pmatrix}$ is an operator that allows the extraction of the linear part, and $J_{1}$, $J_{2}$ are the kinematic Jacobian of the bodies $1$ and $2$
Here, we present the terms $\frac{\partial J_{c}\bm{v}^{+}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$ and $\frac{\partial J_{c}^{\top}\bm{\lambda}^{*}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$ from Section 3.3 using $c$ to be agnostic of the choice of representation of the contact frame $c$. In practice, we use HPP-FCL to get the placement of $c$ relative to the world given the placement of bodies $1$ (${}^{0}M_{1}(\bm{q})$) and $2$ (${}^{0}M_{2}(\bm{q})$) relative to the world. In general, we can write:
${}^{0}M_{c}(\bm{q})$ $=\text{CD}({}^{0}M_{1}(\bm{q}),{}^{0}M_{2}(\bm{q})),$ (44)
where CD is an acronym for collision detection. We are interested in the derivatives of $J_{c}\bm{v}$ when $J_{1}$ and $J_{2}$ are considered constant because their derivation is already considered in the other terms. In this setting, we rewrite (43) apply to a joint velocity $\bm{v}$ as
$J_{c}\bm{v}$ $=E(\operatorname{Ad}_{{}^{0}M_{c}^{-1}(\bm{q}){}^{0}M_{1}(\bm{q})}J_{1}\bm{v}-\operatorname{Ad}_{{}^{0}M_{c}^{-1}(\bm{q}){}^{0}M_{2}(\bm{q})}J_{2}\bm{v}),$ (45)
where $\operatorname{Ad}$ denotes the adjoint operator on $\mathrm{SE}(3)$ to explicitly show the dependency in the variables. We note that only $\bm{q}$ appears in (45).
First, we have
$\frac{\text{d}{}^{0}M_{c}}{\text{d}\theta}$ $=\frac{\partial\text{CD}}{\partial{}^{0}M_{1}}J_{1}\frac{\text{d}\bm{q}}{\text{d}\theta}+\frac{\partial\text{CD}}{\partial{}^{0}M_{2}}J_{2}\frac{\text{d}\bm{q}}{\text{d}\theta},$ (46)
which can be computed using the randomized smoothed derivatives presented in Differentiable collision detection: a randomized smoothing approach [16] in the main paper.
Second, with the rules of spatial algebra, we have for a vector $x$ of the lie algebra, and placement $M$, $M_{a}$, $M_{b}$ in $\mathrm{SE}(3)$
$\text{d}(\operatorname{Ad}_{M}x)$ $=\operatorname{Ad}_{M}\operatorname{ad}_{\text{d}M}x=-\operatorname{ad}_{\operatorname{Ad}_{M}x}\operatorname{Ad}_{M}\text{d}M$ (47a)
$\text{d}(M_{a}^{-1}M_{b})$ $=-\operatorname{Ad}_{M_{b}^{-1}M_{a}}\text{d}M_{a}+\text{d}M_{b},$ (47b)
@@
with $\operatorname{ad}$ the small adjoint on the Lie algebra and by the chain rule
$\text{d}(\operatorname{Ad}_{M_{a}^{-1}M_{b}}x)$ $=\operatorname{ad}_{\operatorname{Ad}_{M_{a}^{-1}M_{b}}x}\text{d}M_{a}-\operatorname{Ad}_{M_{a}^{-1}M_{b}}\operatorname{ad}_{x}\text{d}M_{b},$ (48)
so we have
$\text{d}(J_{c}\bm{v})$ $=E(\operatorname{ad}_{\operatorname{Ad}_{{}^{c}M_{1}}J_{1}\bm{v}}\text{d}{}^{0}M_{c}-\operatorname{Ad}_{{}^{c}M_{1}}\operatorname{ad}_{J_{1}\bm{v}}J_{1}d\bm{q}-\operatorname{ad}_{\operatorname{Ad}_{{}^{c}M_{2}}J_{2}\bm{v}}\text{d}{}^{0}M_{c}+\operatorname{Ad}_{{}^{c}M_{2}}\operatorname{ad}_{J_{2}\bm{v}}J_{2}d\bm{q})$
$=E(\operatorname{ad}_{J_{c}\bm{v}}\text{d}{}^{0}M_{c}-\operatorname{Ad}_{{}^{c}M_{1}}\operatorname{ad}_{J_{1}\bm{v}}J_{1}d\bm{q}+\operatorname{Ad}_{{}^{c}M_{2}}\operatorname{ad}_{J_{2}\bm{v}}J_{2}d\bm{q}),$ (49)
using the linearity of $\operatorname{ad}$ in the index variable. And finally, we obtain the term {mdframed}
$\frac{\partial J_{c}\bm{v}^{+}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$ $=E\left[\left(\operatorname{ad}_{J_{c}\bm{v}^{+}}\frac{\partial\text{CD}}{\partial{}^{0}M_{1}}-{}^{c}X_{1}\operatorname{ad}_{J_{1}\bm{v}^{+}}\right)J_{1}-\left(\operatorname{ad}_{J_{c}\bm{v}^{+}}\frac{\partial\text{CD}}{\partial{}^{0}M_{2}}-{}^{c}X_{2}\operatorname{ad}_{J_{2}\bm{v}^{+}}\right)J_{2}\right]\frac{\text{d}\bm{q}}{\text{d}\theta}.$
@@
and because it is true for any $\bm{v}$ we have $\partial(J_{c}\bm{v})\text{d}\bm{q}=L^{\top}\bm{\lambda}$. We calculate:
$\partial(J_{c}\bm{v})\text{d}\bm{q}$ $=E\left[\left(\operatorname{ad}_{J_{c}\bm{v}}\frac{\partial\text{CD}}{\partial{}^{0}M_{1}}-{}^{c}X_{1}\operatorname{ad}_{J_{1}\bm{v}}\right)J_{1}-\left(\operatorname{ad}_{J_{c}\bm{v}}\frac{\partial\text{CD}}{\partial{}^{0}M_{2}}-{}^{c}X_{2}\operatorname{ad}_{J_{2}\bm{v}}\right)J_{2}\right]\text{d}\bm{q}$
$=L\bm{v}$ $=-E\left[\left(\operatorname{ad}_{\frac{\partial\text{CD}}{\partial{}^{0}M_{1}}J_{1}\text{d}\bm{q}}J_{c}-{}^{c}X_{1}\operatorname{ad}_{J_{1}\text{d}\bm{q}}J_{1}\right)-\left(\operatorname{ad}_{\frac{\partial\text{CD}}{\partial{}^{0}M_{2}}\text{d}\bm{q}}J_{c}-{}^{c}X_{2}\operatorname{ad}_{J_{2}\text{d}\bm{q}}J_{2}\right)\right]\bm{v},$ (51)
using the anti-commutativity of $\operatorname{ad}$. Then we have
$L^{\top}\bm{\lambda}$ $=-\left[\left(J_{c}^{\top}\operatorname{ad}_{\frac{\partial\text{CD}}{\partial{}^{0}M_{1}}J_{1}\text{d}\bm{q}}^{\top}-J_{1}^{\top}\operatorname{ad}_{J_{1}\text{d}\bm{q}}{}^{c}X_{1}^{\top}\right)-\left(J_{c}^{\top}\operatorname{ad}_{\frac{\partial\text{CD}}{\partial{}^{0}M_{2}}J_{2}\text{d}\bm{q}}^{\top}-J_{2}^{\top}\operatorname{ad}_{J_{2}\text{d}\bm{q}}{}^{c}X_{2}^{\top}\right)\right]E\bm{\lambda}$
$=-\left[\left(J_{c}^{\top}\operatorname{P}_{E\bm{\lambda}}\frac{\partial\text{CD}}{\partial{}^{0}M_{1}}J_{1}-J_{1}^{\top}\operatorname{P}_{{}^{c}X_{1}^{\top}E\bm{\lambda}}J_{1}\right)-\left(J_{c}^{\top}\operatorname{P}_{E\bm{\lambda}}\frac{\partial\text{CD}}{\partial{}^{0}M_{2}}J_{2}-J_{2}^{\top}\operatorname{P}_{{}^{c}X_{2}^{\top}E\bm{\lambda}}J_{2}\right)\right]\text{d}\bm{q},$ (52)
where $P$ is the variable commutation of $\operatorname{ad}^{\top}$. Precisely for all $x$ in the Lie algebra and $y$ in the dual Lie algebra: $\operatorname{ad}^{\top}_{x}y=\operatorname{P}_{y}x$. For elements of dual spatial algebra $y=[f,m]$ we have $\operatorname{P}_{y}=\begin{pmatrix}0&f_{\times}\\ f_{\times}&m_{\times}\end{pmatrix}$. And finally, we obtain the second term {mdframed}
@@
$\frac{\partial J_{c}^{\top}\bm{\lambda}^{*}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$ $=\left[\left(J_{c}^{\top}\operatorname{P}_{E\bm{\lambda}^{*}}\frac{\partial\text{CD}}{\partial{}^{0}M_{2}}-J_{2}^{\top}\operatorname{P}_{{}^{c}X_{2}^{\top}E\bm{\lambda}^{*}}\right)J_{2}-\left(J_{c}^{\top}\operatorname{P}_{E\bm{\lambda}^{*}}\frac{\partial\text{CD}}{\partial{}^{0}M_{1}}-J_{1}^{\top}\operatorname{P}_{{}^{c}X_{1}^{\top}E\bm{\lambda}^{*}}\right)J_{1}\right]\frac{\text{d}\bm{q}}{\text{d}\theta}.$
@@
Details on implementation. Here, the terms are calculated for one contact. For multiple contacts, $\frac{\partial J_{c}\bm{v}^{+}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$ is the concatenation of the terms for individual contacts and $\frac{\partial J_{c}^{\top}\bm{\lambda}^{*}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$ is the sum of the terms from each contacts. Note also that similarly to the kinematic Jacobians, the two terms can be computed efficiently by exploiting the sparsity induced by the kinematic structure.
Appendix D - Complete simulation gradients expression
The complete expression of the simulation step derivative is {mdframed}
@@
$\left.\frac{\text{d}\bm{v}^{+}}{\text{d}\theta}\right|_{\lambda=\lambda^{*}}$ $=\frac{\text{d}\bm{v}}{\text{d}\theta}+\Delta t\left(\frac{\partial\text{ABA}}{\partial\bm{q}}\frac{\text{d}\bm{q}}{\text{d}\theta}+\frac{\partial\text{ABA}}{\partial\bm{v}}\frac{\text{d}\bm{v}}{\text{d}\theta}+\frac{\partial\text{ABA}}{\partial\bm{\tau}}\frac{\text{d}\bm{\tau}}{\text{d}\theta}+M^{-1}\frac{\partial J_{c}^{\top}\bm{\lambda}^{*}}{\partial c}\frac{\text{d}c}{\text{d}\theta}\right),$ (53)
@@
Details on implementation. The partial derivatives $\frac{\partial\text{ABA}}{\partial\bm{q},\bm{v},\bm{\tau}}$, $\frac{\partial\text{FKV}}{\partial\bm{q}}$ and the term $M^{-1}(\bm{q})J_{c}^{\top}(\bm{q})$ can be efficiently computed via rigid-body algorithm as implemented in Pinocchio.
The terms $\frac{\partial J_{c}\bm{v}^{+}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$, $\frac{\partial J_{c}^{\top}\bm{\lambda}^{*}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$, $A$, $B$ and $C$ can be computed efficiently jointly with the ABA derivatives during forward and backward search of the kinematic tree to exploit its sparsity.
@@
Therefore, using a Baumgarte correction affects the derivative of the simulation. In particular, the derivatives of the proportional term involve $\Phi$ and thus should be handled when computing the derivatives of the collision detection. Differentiating the derivative term $K_{d}J_{c}v$ is done similarly to the $J_{c}v^{+}$ term i.e. via the Forward Kinematics derivatives. In more details, the term in parentheses of (Appendix D - Complete simulation gradients expression) becomes
@@
Appendix E - Additional experimental support

### v2

@@ Appendix B Implicit NCP gradient system
Appendix C Collision detection contribution
Here, we present the terms $\frac{\partial J_{c}\bm{v}^{+}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$ and $\frac{\partial J_{c}^{\top}\bm{\lambda}^{*}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$ from Section III-D. Given a contact frame $c$ between body $1$ and body $2$, the contact Jacobian is
@@ Appendix C Collision detection contribution
where $E=\begin{pmatrix}\text{Id}_{3}&0_{3,3}\\ 0_{3,3}&0_{3,3}\end{pmatrix}$ is an operator that allows the extraction of the linear part, and $J_{1}$, $J_{2}$ are the kinematic Jacobian of the bodies $1$ and $2$. In practice, the contact placement $c$ is calculated relative the world frame using HPP-FCL. The contact placement is given in function of the placement of two bodies placement relative to the world: ${}^{0}M_{1}(\bm{q})$ and ${}^{0}M_{2}(\bm{q})$. We can write:
${}^{0}M_{c}(\bm{q})$ $={}^{0}\text{CD}({}^{0}M_{1}(\bm{q}),{}^{0}M_{2}(\bm{q})),$ (46)
where ${}^{0}\text{CD}$ is an acronym for collision detection. We are interested in the derivatives of $J_{c}\bm{v}$ when $J_{1}$ and $J_{2}$ are considered constant because their derivation is already considered in the other terms. For clarity of the presentation, we present the calculation for ${}^{c}X_{1}J_{1}v$ as the derivation for $J_{c}v$ follows naturally. In term of Lie groups, ${}^{c}X_{1}=\operatorname{Ad}_{{}^{0}M_{c}^{-1}(\bm{q}){}^{0}M_{1}(\bm{q})}$ where $\operatorname{Ad}$ denotes the adjoint operator on $\mathrm{SE}(3)$ to explicitly show the dependency in the variables. With the rules of spatial algebra, we have for a vector $x$ of the lie algebra, and placement $M$, $M^{\prime}$ in $\mathrm{SE}(3)$
$\text{d}(\operatorname{Ad}_{M}x)$ $=-\operatorname{ad}_{\operatorname{Ad}_{M}x}\operatorname{Ad}_{M}\text{d}M$ (47a)
$\text{d}(M^{-1}M^{\prime})$ $=-\operatorname{Ad}_{M^{\prime-1}M}\text{d}M+\text{d}M^{\prime},$ (47b)
@@
with $\operatorname{ad}$ the small adjoint on the Lie algebra. By the chain rule we obtain
$\text{d}(\operatorname{Ad}_{M^{-1}M^{\prime}}x)$ $=\operatorname{ad}_{\operatorname{Ad}_{M^{-1}M^{\prime}}x}\text{d}M-\operatorname{Ad}_{M^{-1}M^{\prime}}\operatorname{ad}_{x}\text{d}M^{\prime}.$ (48)
Applied to $M={}^{0}M_{c}(\bm{q})$ and $M^{\prime}={}^{0}M_{1}(\bm{q})$ we obtain
$\text{d}({}^{c}X_{1}(\bm{q})J_{1}v)$ $=\operatorname{ad}_{{}^{c}X_{1}J_{1}\bm{v}}\text{d}{}^{0}M_{c}-{}^{c}X_{1}\operatorname{ad}_{J_{1}\bm{v}}\text{d}{}^{0}M_{1}$ (49)
$=({}^{c}X_{1}J_{1}\bm{v})\times\text{d}{}^{0}M_{c}-{}^{c}X_{1}(J_{1}\bm{v}\times\text{d}{}^{0}M_{1}).$ (50)
Where in the second form we use the generalized cross product from the notation of Featherstone [16]. The calculus is similar for $J_{2}$. Now differentiating the collision detection, we have
$\frac{\text{d}{}^{0}M_{c}}{\text{d}\theta}$ $=\frac{\partial{}^{0}\text{CD}}{\partial{}^{0}M_{1}}J_{1}\frac{\text{d}\bm{q}}{\text{d}\theta}+\frac{\partial{}^{0}\text{CD}}{\partial{}^{0}M_{2}}J_{2}\frac{\text{d}\bm{q}}{\text{d}\theta},$ (51)
which can be computed using the randomized smoothed derivatives presented in Differentiable collision detection: a randomized smoothing approach [16] in the main paper. For self explanation of the paper, we provide the final formula: {mdframed}
$\frac{\partial J_{c}\bm{v}^{+}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$ $=E\left[\left(({J_{c}\bm{v}^{+}})_{\times}\frac{\partial{}^{0}\text{CD}}{\partial{}^{0}M_{1}}-({}^{c}X_{1}{J_{1}\bm{v}^{+}})_{\times}\right)J_{1}+\left(({J_{c}\bm{v}^{+}})_{\times}\frac{\partial{}^{0}\text{CD}}{\partial{}^{0}M_{2}}+({}^{c}X_{2}{J_{2}\bm{v}^{+}})_{\times}\right)J_{2}\right]\frac{\text{d}\bm{q}}{\text{d}\theta}.$
@@
and because it is true for any $\bm{v}$ we have $\partial(J_{c}^{\top}\bm{\lambda})\text{d}\bm{q}=L^{\top}\bm{\lambda}$. We calculate $L$ using the anti-commutativity of $\operatorname{ad}$:
$\partial(J_{c}\bm{v})\text{d}\bm{q}$ $=E\left[\left(\operatorname{ad}_{J_{c}\bm{v}}\frac{\partial{}^{0}\text{CD}}{\partial{}^{0}M_{1}}-{}^{c}X_{1}\operatorname{ad}_{J_{1}\bm{v}}\right)J_{1}+\left(\operatorname{ad}_{J_{c}\bm{v}}\frac{\partial{}^{0}\text{CD}}{\partial{}^{0}M_{2}}+{}^{c}X_{2}\operatorname{ad}_{J_{2}\bm{v}}\right)J_{2}\right]\text{d}\bm{q}$
$L\bm{v}$ $=-E\left[\left(\operatorname{ad}_{\frac{\partial{}^{0}\text{CD}}{\partial{}^{0}M_{1}}J_{1}\text{d}\bm{q}}J_{c}-{}^{c}X_{1}\operatorname{ad}_{J_{1}\text{d}\bm{q}}J_{1}\right)+\left(\operatorname{ad}_{\frac{\partial{}^{0}\text{CD}}{\partial{}^{0}M_{2}}\text{d}\bm{q}}J_{c}+{}^{c}X_{2}\operatorname{ad}_{J_{2}\text{d}\bm{q}}J_{2}\right)\right]\bm{v}.$ (53)
Taking the transpose and introducing the operator $P$ is the variable commutation of $\operatorname{ad}^{\top}$. Precisely for all $x$ in the Lie algebra and $y$ in the dual Lie algebra: $\operatorname{ad}^{\top}_{x}y=\operatorname{P}_{y}x$ we obtain:
$L^{\top}\bm{\lambda}$ $=-\left[\left(J_{c}^{\top}\operatorname{ad}_{\frac{\partial{}^{0}\text{CD}}{\partial{}^{0}M_{1}}J_{1}\text{d}\bm{q}}^{\top}-J_{1}^{\top}\operatorname{ad}^{\top}_{J_{1}\text{d}\bm{q}}{}^{c}X_{1}^{\top}\right)+\left(J_{c}^{\top}\operatorname{ad}_{\frac{\partial{}^{0}\text{CD}}{\partial{}^{0}M_{2}}J_{2}\text{d}\bm{q}}^{\top}+J_{2}^{\top}\operatorname{ad}^{\top}_{J_{2}\text{d}\bm{q}}{}^{c}X_{2}^{\top}\right)\right]E\bm{\lambda}$
$\partial(J_{c}^{\top}\bm{\lambda})\text{d}\bm{q}$ $=-\left[\left(J_{c}^{\top}\operatorname{P}_{E\bm{\lambda}}\frac{\partial{}^{0}\text{CD}}{\partial{}^{0}M_{1}}J_{1}-J_{1}^{\top}\operatorname{P}_{{}^{c}X_{1}^{\top}E\bm{\lambda}}J_{1}\right)+\left(J_{c}^{\top}\operatorname{P}_{E\bm{\lambda}}\frac{\partial{}^{0}\text{CD}}{\partial{}^{0}M_{2}}J_{2}+J_{2}^{\top}\operatorname{P}_{{}^{c}X_{2}^{\top}E\bm{\lambda}}J_{2}\right)\right]\text{d}\bm{q}.$ (54)
And finally, we obtain the second term {mdframed}
@@
$\frac{\partial J_{c}^{\top}\bm{\lambda}^{*}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$ $=\left[\left(J_{1}^{\top}\operatorname{P}_{{}^{c}X_{1}^{\top}E\bm{\lambda}^{*}}-J_{c}^{\top}\operatorname{P}_{E\bm{\lambda}^{*}}\frac{\partial{}^{0}\text{CD}}{\partial{}^{0}M_{1}}\right)J_{1}-\left(J_{2}^{\top}\operatorname{P}_{{}^{c}X_{2}^{\top}E\bm{\lambda}^{*}}+J_{c}^{\top}\operatorname{P}_{E\bm{\lambda}^{*}}\frac{\partial{}^{0}\text{CD}}{\partial{}^{0}M_{2}}\right)J_{2}\right]\frac{\text{d}\bm{q}}{\text{d}\theta}.$
@@
Details on implementation. Here, the terms are calculated for one contact. For multiple contacts, $\frac{\partial J_{c}\bm{v}^{+}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$ is the concatenation of the terms for individual contacts and $\frac{\partial J_{c}^{\top}\bm{\lambda}^{*}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$ is the sum of the terms from each contacts. Note that for elements of dual spatial algebra $y=[f,m]$ we have the closed form $\operatorname{P}_{y}=\begin{pmatrix}0&f_{\times}\\ f_{\times}&m_{\times}\end{pmatrix}$. Note also that similarly to the kinematic Jacobians, the two terms can be computed efficiently by exploiting the sparsity induced by the kinematic structure.
Appendix D Complete simulation gradients expression
Combining the terms from the collision detection, the term from the velocity forward kinematic, and the main calculation presented in Section III, the complete expression of the simulation step derivative is {mdframed}
@@ Appendix D Complete simulation gradients expression
$\left.\frac{\text{d}\bm{v}^{+}}{\text{d}\theta}\right|_{\lambda=\lambda^{*}}$ $=\frac{\text{d}\bm{v}}{\text{d}\theta}+\Delta t\left(\frac{\partial\text{UFD}}{\partial\bm{q}}\frac{\text{d}\bm{q}}{\text{d}\theta}+\frac{\partial\text{UFD}}{\partial\bm{v}}\frac{\text{d}\bm{v}}{\text{d}\theta}+\frac{\partial\text{UFD}}{\partial\bm{\tau}}\frac{\text{d}\bm{\tau}}{\text{d}\theta}+M^{-1}\frac{\partial J_{c}^{\top}\bm{\lambda}^{*}}{\partial c}\frac{\text{d}c}{\text{d}\theta}\right),$ (55)
@@
Details on implementation. The partial derivatives $\frac{\partial\text{UFD}}{\partial\bm{q},\bm{v},\bm{\tau}}$, $\frac{\partial\text{FKV}}{\partial\bm{q}}$ and the term $M^{-1}(\bm{q})J_{c}^{\top}(\bm{q})$ can be efficiently computed via rigid-body algorithm as implemented in Pinocchio. The terms $\frac{\partial J_{c}\bm{v}^{+}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$, $\frac{\partial J_{c}^{\top}\bm{\lambda}^{*}}{\partial c}\frac{\text{d}c}{\text{d}\theta}$, $A$, $B$ and $C$ can be computed efficiently jointly with the ABA derivatives during forward and backward search of the kinematic tree to exploit its sparsity.
@@
Therefore, using a Baumgarte correction affects the derivative of the simulation. In particular, the derivatives of the proportional term involve $\Phi$ and thus should be handled when computing the derivatives of the collision detection. Differentiating the derivative term $K_{d}J_{c}v$ is done similarly to the $J_{c}v^{+}$ term i.e. via the Forward Kinematics derivatives. In more details, the term in parentheses of (D) becomes
@@
Appendix E Additional experimental support